Question #17,18,19, and 20

Question #17,18,19, and 20
Answer #1

17

Using law of conservation of momentum

m1u1 + m2u2 = m1v1 + m2v2

v = 10x5 + 2 x0 / (10+2) = 50/12 = 4.17 m/s

all other are based on same formula